OpenStudy (anonymous):

There are 12 girls and 8 boys in Mrs. Martins class. Each day, she randomly asks one student to take attendance. In 180 school days, which is the best prediction for the number of times that a student will be a girl? * A.72 B.90 C.99 D.108

OpenStudy (anonymous):

\[\frac{12}{20}\] is the ratio of girls to the total multiply \[\frac{12}{20}\times 180\]